Jenne and Koji
We tweaked the alignment of the TT mirror.
First we put a G&H mirror, but the mirror was misaligned and touching the ECD as the magnet was too heavy. We tried to move the wires towards the magnet by 1mm.
It was not enough but once we moved the clamps towards the magnet, we got the range to adjust the pitching back and forth.
We tried to align it by the feaher touch to the clamp, we could not get close to the precision of 10mrad as the final tightening of the clamp screws did change the alignment.
We will try to adjust the fine alignment tomorrow again.
The damping in pitch, yaw and longitudinal looks quite good. We will also try to characterize the damping of the suspension using a simple oplev setup.